2. The accuracy of taking orders at the car service window in fast food places is essential for customer satisfaction. Accuracy is measured as a percentage of orders for a main item, a secondary item and a beverage that are served correctly. Recently, the percentage of pediods at Wendy's was 88%. Assume that you and two of your friends are in your car at Wendy's window and each makes your order as described above.

a. What is the probability that all three orders will be handled accurately?

b. None of the three?

C. At least two of the three?
Answer #1

Binomial distribution: P(x) = nCx px qn-x

P(an accurate order), p = 0.88

q = 1 - 0.88 = 0.12

Sample size, n = 3

a. P(all three orders will be handled accurately), P(3) = 0.883

= 0.6815

b. P(none of the three), P(0) = 0.123

= 0.0017

c. P(at least two of the three), P(2) + P(3) = 3x0.882x0.12 + 0.883

= 0.9603